# Ctrl-Crash: Controllable Diffusion for Realistic Car Crashes

### v1

@@ A.2 Additional Quantitative Results: Ablations
Appendix B FFT-Based Filtering Heuristic
In the context of video dataset curation, it is important to ensure that the frames are of genuine high quality and not artificially upscaled from lower resolutions. When dealing with in-the-wild dashcam footage from various online sources, it is often difficult to guarantee high-quality samples and many videos present compression or resizing artifacts that can dramatically affect the dataset quality. As a first step in the video processing pipeline for this work (see Appendix C) we designed a function to estimate which video samples may be originally of very low resolution using frequency analysis. The function is used as a quality control step to automatically detect and potentially filter out such frames or videos.
We define a "upscaling factor estimation" function which is designed to estimate how much an image (a video frame) has been upscaled before being resized to a standard resolution (e.g., 720×1280). Upscaling refers to the process of enlarging an image, which can introduce artifacts and degrade quality. This function provides a quantitative heuristic score indicating the likelihood and degree of upscaling, which can be used to filter out low-quality or artificially enlarged video samples in a dataset. The present the procedure textually below and more formally in Algorithm 1.
Here is a step-by-step breakdown of the process:
  1. 1.
    Image Loading Load a video frame in grayscale to simplify the analysis by focusing on structural content rather than color information.
  2. 2.
    Fourier Transform Compute the 2D Fast Fourier Transform (FFT) of the image, which converts the spatial domain image into the frequency domain. The result is then shifted so that the low-frequency components are centered.
  3. 3.
    Magnitude Spectrum Calculate the magnitude (absolute value) of the frequency components, representing the strength of different frequency components in the image.
  4. 4.
    High-Frequency Energy Calculation Define a circular region at the center of the frequency spectrum (corresponding to low frequencies) and create a mask to exclude this central region, thereby isolating the high-frequency components. The sum of the magnitudes outside the central region (i.e., the high-frequency energy) is computed.
  5. 5.
    Normalization Normalize the high-frequency energy by the total number of pixels, yielding an energy score that is independent of image size.
  6. 6.
    Upsizing Factor Estimation Compute an "upsizing factor" as the inverse of (1 + energy score). The rationale is that upscaled images tend to have lower high-frequency energy (since upscaling smooths out details), so a lower energy score indicates more upscaling. The formula ensures that the factor decreases as the high-frequency energy decreases.
  7. 7.
    Return Value Return the upsizing factor, which can be interpreted as a proxy for the degree of upscaling: Lower values indicate more likely and/or more severe upscaling. Higher values indicate less or no upscaling (i.e., more genuine high-frequency detail).
@@ Appendix C Video Processing
For training, we use the MM-AU dataset [10], a large-scale collection of dashcam crash videos collected from online sources. To ensure high quality we curate this dataset by applying a series of filtering steps:
@@
    Low-resolution or heavily compressed videos are filtered out using FFT-based heuristics to detect blocky or low-fidelity regions and prioritize videos with clear visibility and identifiable motion patterns for training. See Appendix B for more details.
@@
    PySceneDetect [6] is used to remove videos with shot changes or very sudden unnatural camera angle changes (e.g., dashcam falls over)
@@
    Videos are sampled at a consistent frame rate of 6 fps, and cropped to a resolution of $512\times 320$ (width $\times$ height) to fit our model’s input format and remove unwanted watermarks in the process. For training, we trim the videos into 25-frame clips ($\approx$4s videos) and therefore reject any clips with less than 25 frames.
@@
    For this work, we have opted to filter out all accidents involving visible humans to avoid teaching the model to generate gruesome and violent scenes. This includes all accidents directly involving: pedestrians, cyclists, and motorbikes.
Additionally, many video samples were reviewed manually (reviewing tools available on the project github) to identify low-quality video samples that were not identified by the processing pipeline. After these filtering steps, we retain approximately 7,500 videos from the original 11,727 videos. We split these videos into a training set and a held-out test set by randomly sampling by accident type categories with a 90/10 ratio.
Appendix D Bounding Box Extraction Pipeline
Figure 11: Shows sequence of frames with a car going into the ditch. SAM2’s mask predictions for the two cars in the video are shown in purple and orange. YOLO’s first prediction of the purple colored car is shown in frame 16 with a light pink bounding box. SAM2 works bidirectionally and can therefore infer the appearance of the purple-colored car many frames earlier and until the very last instant before it disappears into the ditch; this would not be possible with a tracker solely based on YOLO which relies on the object to be recognizable as one of the target classes in order to be detected.
We develop a hybrid pipeline, referred to as YOLO-SAM, to extract high-quality bounding box annotations for all visible road users in crash and driving videos. The pipeline combines the fast, class-aware object detection capabilities of YOLOv8 [23] with the instance-level segmentation and temporal consistency of SAM2 (Segment Anything Model v2)[33]. This combination addresses limitations in both detection and tracking, especially under occlusion, deformation, or abrupt motion—common in crash scenarios—and supports dynamic entry and exit of agents
YOLO Detection and Initial Tracking
YOLOv8 is applied to each frame to detect objects and assign initial bounding boxes. This step provides a fast and reliable detection of objects in each frame. YOLO also attempts to track objects across frames. However, it has several limitations, namely:
@@ Appendix D Bounding Box Extraction Pipeline
    Multiple Detections: YOLO may detect the same object multiple times at different time points. To address this, the code checks for overlap between its own predictions and rejects duplicates if the overlap is too high (if IOU > 0.8).
@@
    Loss of Detection: YOLO may lose detection of an object through time, especially in complex scenarios where objects become unrecognizable due to occlusion or deformation. This is where SAM2 comes into play, as it is better at tracking objects over time.
SAM2 Segmentation and Tracking
SAM2 is used to produce instance-level masks for each object initially detected by YOLO. This allows for more precise tracking, especially when objects are occluded or deformed. Concretely, SAM2 is used to handle the following YOLO limitations:
@@
    Distorted Bounding Boxes: YOLO may distort the size of the bounding box. SAM2 tracks the shape more reliably, ensuring that the bounding boxes are accurate.
@@
    Redetection with New Track ID: If YOLO loses detection and then redetects an object with a new track ID, the code checks for overlap between YOLO’s new prediction and SAM’s prediction. If the overlap is too high, YOLO’s new prediction is rejected.
@@
    Track ID Switching: If YOLO assigns an existing track ID to a different object after losing detection, the code changes this track ID to a new value, as it does not trust YOLO for tracking. This assumes that the new detection is actually a new object, which is verified by the overlap check with SAM’s prediction.
@@
    Late Detection: YOLO may detect an object late (i.e., several frames after the object entered the frame). SAM2 works bidirectionally and can help by making predictions in earlier frames based on the bounding box prompt in later frames. However, we have found this to be risky as SAM2 may make mistakes and is biased towards thinking all objects stay in frame during the clip. Therefore, the code only accepts SAM’s "early" predictions a few frames ahead of YOLO’s predictions.
Combining YOLO and SAM2 Predictions
The code filters the predictions from both YOLO and SAM2 to ensure that only reliable bounding boxes are retained. Finally, we retain the SAM2 predictions for each frame and convert the predicted masks to tight bounding boxes. The final output is a list of dictionaries, each representing a frame of the video. Each frame contains a list of tracked objects with their bounding boxes, track IDs, and class names. Code for this implementation is made publicly available on our project page.
@@ Appendix F FVD and JEDi Metrics Computation
We use two distributional metrics to evaluate video generation quality: Fréchet Video Distance (FVD) [41] and JEDi [29]. Both metrics aim to measure the distance between the distribution of generated videos and real videos. FVD is computed by embedding videos into the feature space of a pretrained Inflated 3D ConvNet (I3D), assuming a multivariate Gaussian distribution, and then computing the Fréchet distance. JEDi, in contrast, uses features from a Joint Embedding Predictive Architecture (JEPA) and computes distributional similarity via Maximum Mean Discrepancy (MMD) with a polynomial kernel, avoiding the Gaussian assumption and improving temporal sensitivity and sample efficiency. Both are computed on extracted features from generated and ground-truth video samples.
We adopt two evaluation protocols:
@@
    Unpaired Evaluation (Random Ground-Truth Sampling): For AVD2, which provides 2000 unpaired generated videos (i.e., without known ground-truth counterparts), we compute FVD and JEDi by comparing to a set of 500 randomly sampled videos from the MM-AU validation set. This approach estimates the distance between the overall distribution of generated videos and real crash scenarios, and enables fair comparison across models lacking GT alignment. For consistency, we also use this protocol to evaluate Ctrl-Crash, SVD Base, and Ctrl-V by sampling 200 generated videos from each model. These results are reported in Table 1 of the main paper.
@@
    Paired Evaluation (Ground-Truth Matching): For methods like Ctrl-Crash, Ctrl-V, and SVD Base—where ground-truth video correspondence is available—we compute FVD and JEDi between 200 generated videos and their matched ground-truth counterparts from MM-AU. This approach captures alignment not only in distribution but also in per-sample temporal and visual fidelity. These results are reported in Table 2 and Table 3 of the main paper.
For OAVD and DrivingGen, we report the FVD values cited in their original publications due to the unavailability of code or generated samples. These are marked with an asterisk (“*”) in Table 1 of the main paper.

### v2

@@ A.1 Additional Quantitative Results: Ablations
Table 4: Effect of crash type conditioning on crash video generation quality. We evaluate Ctrl-Crash on the Crash Counterfactuals task by varying the crash type conditioning and nothing else. For each case, 200 videos were generated using the same initial image and three bounding box frames as initial context, but with different desired crash types.
[…]
GT crash type 375.7 0.2949
0 - no crash 400.9 0.4514
1 - ego-only 379.5 0.3091
2 - ego/vehicle 372.9 0.3001
3 - vehicle-only 398.1 0.3856
4 - vehicle/vehicle 383.4 0.3416
In Table 4, we evaluate Ctrl-Crash on the Crash Counterfactuals task by varying only the crash type while keeping the initial image and partial bounding box sequence fixed. The model achieves its best performance on ego/vehicle crashes (type 2), likely due to their higher frequency in the training data, while vehicle-only crashes (type 3) are harder to generate plausibly, reflecting their visual complexity. Interestingly, ego-only crashes (type 1), though underrepresented in the dataset, also perform well—possibly because they involve simpler dynamics without visible impact. Overall, counterfactual generations remain close in quality to the ground-truth-guided baseline, suggesting that Ctrl-Crash can produce realistic alternate outcomes, though performance declines when strong visual cues in the initial frame conflict with the intended crash type. Class frequency in the training set are provided in Appendix H.
A.2 Additional Qualitative Results
SOTA models
Figure 6: State-of-the-art text-to-video diffusion models struggle to generate a realistic car accident (as shown in main paper). Top: Nvidia Cosmos (7B Text2World), Middle: OpenAI Sora, Bottom: DeepMind Veo3.
State-of-the-art diffusion-based video models, such as OpenAI’s Sora [34], Nvidia’s Cosmos [19], and DeepMind’s Veo3 [35] fail to generate plausible crash scenarios. We tested these three models with several samples and though they tend to generate high definition and smooth video, they consistently fail to generate physically plausible crashes (if any crash at all), we show some examples in Figure 6.
In the top row of the figure we show a sample result from Nvidia Cosmos-Predict1-7B-Text2World that was instructed with the following text prompt: ”On a highway two cars collide at very fast speeds head-on”. This produces a highly implausible scene, where the car on the left suddenly starts to levitate from its rear-end and a cloud of smoke resembling an explosion appears, followed by disjointed fragments of torn metal emerging from the ground that transforms in a dark vehicle rushing towards the camera.
The Nvidia Cosmos model suite has been specifically trained for physics-aware generation, but still fail to generate a car crash. Alternatively, we show a sample from OpenAI’s Sora video model in Figure 6 in the middle row. The Sora video model was given the prompt: ”At an intersection two cars collide with each other at full speed resulting in a crash”. The resulting video shows a car that spins erratically, changes shape and direction, and produces visible artifacts—culminating in a pile of twisted metal and glass.
Finally, we show an example using DeepMind’s Veo3 in the bottom row of Figure 6. We used the following prompt: ”At an intersection two cars collide with each other at full speed resulting in a crash”. Perhaps the closest to depicting a car crash when looking at individual frames, but when watching the video as a whole it becomes clear that the events are completely unrealistic. In the generated video, we see two still vehicles very close to each other that start accelerating, then we see debris and smoke as the cars slowly move towards each other, then the front of the cars phase through each other and reveal completely undamaged cars, and finally the red car drives away from the scene.
In contrast, through our approach of sourcing crash data, preprocessing and annotating sequences with bounding boxes, crash type conditioning, and our stochastic conditioning approach, Ctrl-Crash can much more reliably generate physically plausible crash imagery and collision dynamics.
Figure 7: Crash scenarios generated by Ctrl-Crash from non-crash scenes in the BDD100k dataset. Each row shows 5 frames from a generated 25-frame clip. Despite originating from benign driving scenes, the model produces visually coherent and diverse crash outcomes across examples.
Generating Crashes from Non-Crash Data
To demonstrate the practical utility of Ctrl-Crash for safety-critical dataset augmentation, we apply our model to generate crash scenarios seeded from the BDD100k dataset, an established real-world driving dataset that contains no actual crashes. By conditioning on initial frames and coarse agent trajectories from BDD100k, Ctrl-Crash is able to hallucinate plausible accident outcomes, transforming otherwise uneventful driving clips into diverse crash scenarios. This capability is especially valuable for enriching datasets used in autonomous driving research, where real crash data is scarce, sensitive, or ethically challenging to collect. Synthesizing rare and hazardous events from benign scenes enables safer training and evaluation of perception and planning systems without requiring exposure to real-world danger. In Figure 7, we show some examples of car crash videos generated by Ctrl-Crash from non-crash scenes in the BDD100k dataset.
Car Crash Video Diffusion Models Comparison
Other recent work [24, 25] focus on the task of car crash video generation, but lack the visual quality for convincing car crashes. As mentioned in the main paper (and presented in Appendix I), we conducted a user survey that shows that Ctrl-Crash significantly outperforms these existing methods for visual fidelity and physical plausibility of generated car crashes. We show additional qualitative comparisons in Figure 8 and Figure 9, along with many samples generated from Ctrl-Crash in Figure 10, and Figure 11.
Figure 8: Qualitative comparison of ”rear-end crashes” between different methods. For each method, we show 5 frames from the video along with either a green check mark if there appears to have a crash in the video otherwise a red ’X’. From top to bottom: SVD (stable-video-diffusion-img2vid) [9] prompted with the initial frame from a rear-end crash video, we see some normal driving but very inconsistent lighting and color shades with visible distorsions. AVD2 [25], we see what appears to be a rear-end crash with a very distorted leading vehicle and background. DrivingGen [24], we see a rear-end crash with a leading vehicle that changes appearance every frame. Overall the video is very choppy with little temporal consistency. Cosmos (Cosmos-Predict1-7B-Video2World) [19], prompted with text suggesting a rear-end crash and 9 initial images where a car is rapidly approaching a truck, the predicted frames show the car unrealistically shrinking as it approaches the truck without any signs of a collision. Ctrl-V [6], prompted with a sequence of bounding-boxes suggesting a rear-end crash with a leading car, we see the leading car keep its distance and not crash occurs. Ctrl-Crash (ours): prompted with the same bounding box sequence as Ctrl-V and with the discrete crash type ”ego/vehicle crash”, we see a physically plausible rear-end collision with the ego vehicle visibly shaking from the impact.
Figure 9: Qualitative comparison of ”t-bone crashes” between different methods. For each method, we show 6 frames from the video along with either a green check mark if there appears to have a crash in the video otherwise a red ’X’. From top to bottom: SVD (stable-video-diffusion-img2vid) [9] prompted with the initial frame from a t-bone car crash video, we see blurry vehicle a distorted motion blur as it drives in front of the ego vehicle without any collision. AVD2 [25], we can make out what seems to be a t-bone crash with a heavily distorted black car. There are many artifacts and temporal inconsistencies which makes the sequence of events hard to follow. DrivingGen [24], a gray sedan drive in front of the ego vehicle progressively getting closer until it seems to collide with it. Motion is jerky and uneven between timesteps and the appearance of the gray car shapes almost every frame. Cosmos (Cosmos-Predict1-5B-Video2World) [19], prompted with creating a t-bone crash and 9 initial frames showing a car turn in front of the ego car, we see the leading car start to distort as the ego approaches it and then it shrivels and shrinks until it nearly disappears. Ctrl-V [6], prompted with a sequence of bounding-boxes suggesting a t-bone crash with a car incoming from the left, we see a car drive in from the left and then just passed the ego car without any collision. Ctrl-Crash (ours): prompted with the same bounding box sequence as Ctrl-V and with the discrete crash type ”ego/vehicle crash”, we see a physically plausible t-bone collision with the a black sedan incoming from the left.
Figure 10: Ctrl-Crash qualitative results conditioned on an initial 9 bounding box frames (i.e., the first two frames of these sequences were conditioned on bounding box frames, but not the others).
Figure 11: Ctrl-Crash qualitative results conditioned on 25 (all) bounding box frames.
Appendix B Low Resolution Filtering Heuristic
@@
2:  Load image in grayscale: $I\leftarrow\texttt{cv2.imread(image\_path, GRAYSCALE)}$
@@
A key challenge in curating video datasets from online sources is ensuring that the selected samples reflect genuine high-resolution content rather than upscaled of heavily compressed footage (via resizing). This is particularly relevant for dashcam videos, where visual clarity and high-frequency details are critical for learning accurate dynamics and semantics. As part of the video preprocessing pipeline (see Section C), we introduce a frequency-based filtering heuristic to estimate the degree of upscaling in video frames.
The heuristic is based on the observation that upscaled or low-quality images tend to exhibit diminished high-frequency content due to interpolation artifacts and smoothing. By analyzing the energy distribution in the frequency domain, we define an upsizing factor that acts as a proxy for the likelihood of artificial upscaling. The method proceeds as follows:
  1. 1.
    Image Loading. Each video frame is converted to grayscale to simplify analysis and focus on structural image content rather than color channels.
  2. 2.
    Fourier Transform. A 2D Fast Fourier Transform (FFT) is applied to the grayscale image to obtain its frequency representation. The spectrum is shifted so that low-frequency components are centered.
  3. 3.
    Magnitude Spectrum. The magnitude of the frequency components is computed by taking the absolute value of the FFT coefficients. This magnitude spectrum represents the intensity of different spatial frequencies in the image.
  4. 4.
    High-Frequency Energy Calculation. A circular mask is applied to exclude the low-frequency region in the center of the spectrum. The sum of magnitudes outside this central region defines the high-frequency energy.
  5. 5.
    Normalization. The high-frequency energy is normalized by the total number of pixels to produce a scale-invariant energy score.
  6. 6.
    Upsizing Factor Estimation. The final upsizing factor $U$ is computed as: $U=\frac{1}{1+E}$, where $E$ is the normalized high-frequency energy. This formulation ensures that frames with low high-frequency energy—indicative of potential upscaling—yield lower scores.
  7. 7.
    Interpretation. The resulting upsizing factor provides a heuristic score: lower values indicate a higher likelihood of artificial enlargement or blur, while higher values suggest more genuine, detailed high-resolution frames.
This frequency-domain metric serves as a lightweight yet effective automated quality check, helping to filter out training samples that lack meaningful visual detail. The corresponding implementation is summarized formally in Algorithm 1.
@@ Appendix C Video Processing
To train our model, we use the MM-AU dataset [11], a large-scale collection of in-the-wild dashcam crash videos sourced from public platforms. However, due to the variable quality and content of these videos, we curate the dataset using a multi-stage filtering and preprocessing pipeline. The process can be summarized as follows:
@@
    Filtering out low-quality, compressed, or blocky videos based on the low-resolution heuristics (described in Appendix B).
@@
    Removing clips with shot changes or unnatural scene cuts using scene detection.
@@
    Normalizing the format: frame rate, resolution, clip length.
@@
    Excluding scenes involving visible humans (e.g., pedestrians, cyclists, motorbikes).
In the first step, we apply frequency-domain heuristics to identify and remove videos that suffer from excessive compression artifacts or poor resolution. These issues can hinder the model’s ability to learn coherent motion and object dynamics. We use a Fast Fourier Transform (FFT)-based method to detect videos with large blocky regions or low-frequency dominance—signals of strong compression or blur. This approach, described in Appendix B, helps prioritize clips with high visual clarity and well-defined agent motion.
Next, we use PySceneDetect [36] to identify and discard videos containing abrupt scene transitions or camera disruptions—such as sudden viewpoint changes or dashcams that fall mid-recording. Removing such discontinuities ensures more stable temporal consistency within each training clip.
We then normalize all video clips to a consistent resolution of $512\times 320$ (width × height) and sample them at 6 frames per second. This not only aligns with the spatial and temporal input requirements of our model but also helps crop out unwanted overlays such as watermarks that often appear near video borders. We segment each video into fixed-length clips of 25 frames (approximately 4 seconds), discarding any segments shorter than this threshold.
Finally, for ethical and safety considerations, we exclude all videos depicting visible humans involved in crashes. This includes scenes involving pedestrians, cyclists, or motorcyclists. Our intention is to avoid training the model to depict harmful or distressing scenarios involving vulnerable road users.
In addition to the automated filtering steps, we also perform manual review of many video samples using internal tooling (made available on the project GitHub). This step helps identify edge cases and poor-quality examples that may have bypassed the automated pipeline, ensuring a cleaner and more appropriate training set.
Appendix D YOLO-SAM: Hybrid Bounding Box Annotation Pipeline
Figure 12: Example of bounding box tracking for a vehicle veering off the road. SAM2 segmentations for the two vehicles are shown in purple and orange. YOLOv8 detects the purple car only starting at frame 16 (light pink box), while SAM2 successfully infers its presence across many more frames by propagating the segmentation bidirectionally in time. This provides more complete and temporally consistent annotations, especially in cases where objects become partially occluded or deformed.
We develop a hybrid annotation pipeline, YOLO-SAM, to generate high-quality bounding boxes for visible road users in crash and driving videos. This combines the speed and class-awareness of YOLOv8 [27] with the instance-level segmentation and temporal propagation capabilities of SAM2 [28]. The hybrid design addresses detection failures due to occlusion, deformation, or sudden motion—common in crash scenarios—and supports dynamic entry/exit of agents.
YOLOv8 is run frame-by-frame to detect standard object classes (e.g., cars, trucks, buses) and assign initial bounding boxes and track IDs. While fast and accurate under normal conditions, YOLO has several shortcomings:
@@ Appendix D YOLO-SAM: Hybrid Bounding Box Annotation Pipeline
    Duplicate Detections: YOLO can assign multiple detections to the same object over time. We reject predictions if their IoU with previous detections exceeds 0.8.
@@
    Tracking Loss: YOLO may lose an object (e.g., under occlusion) and redetect it with a new track ID. These failures are corrected using SAM2.
To address YOLO’s limitations, we use SAM2 to refine and extend YOLO detections:
@@
    Shape Correction: SAM2 provides temporally consistent masks, reducing YOLO’s tendency to distort box sizes and shapes.
@@
    Redetection Verification: When YOLO redetects a lost object with a new ID, we compare its spatial overlap with SAM2’s mask to either restore the original ID or assign a new one.
@@
    Track ID Switches: If YOLO mistakenly assigns an existing ID to a new object, SAM2’s predictions are used to detect the mismatch and correct the ID.
@@
    Early-frame Recovery: SAM2 supports bidirectional propagation, allowing recovery of early frames missed by YOLO. However, to avoid hallucinated presence, we only accept early-frame completions up to a few frames before YOLO’s first detection, and only when confidence is high.
Figure 12 shows an example where a vehicle exits the road and YOLO detects it only mid-sequence, while SAM2 successfully tracks it throughout by propagating the mask temporally.
For each YOLO-detected object, SAM2 is prompted to generate per-frame instance masks. These masks are converted to tight bounding boxes, and tracking continuity is enforced by ID consistency checks based on IoU overlap. The result is a per-frame annotation of bounding boxes with consistent track IDs and object classes.
Compared to YOLO or SAM2 alone, the hybrid approach provides significantly more complete and robust annotations, crucial for training controllable video generation models like Ctrl-Crash. The full implementation is available in our project repository for reproducibility.
@@ Appendix F FVD and JEDi Metrics Computation
We evaluate the quality of generated videos using both distributional video-level metrics and frame-level fidelity metrics. Fréchet Video Distance (FVD) [31] measures the distance between the distributions of generated and real videos (lower is better). This is done by embedding the two distributions in the feature space of a pretrained Inflated 3D ConvNet (I3D) and computing the Fréchet distance under the assumption of a multivariate Gaussian. JEDi [32] addresses limitations of FVD by relaxing the Gaussian assumption, employing features from a Joint Embedding Predictive Architecture (JEPA), and using Maximum Mean Discrepancy (MMD) with a polynomial kernel to improve both temporal sensitivity and sample efficiency. Together, these metrics provide complementary views of video realism at the distributional level.
To evaluate distributional similarity between generated and real videos, we compute Fréchet Video Distance (FVD) and JEDi under two complementary evaluation protocols, which differ based on how the ground-truth distribution is constructed:
@@
    Condition-Aligned Evaluation: For models such as Ctrl-Crash, Ctrl-V, and SVD Base, we have access to the ground-truth data used to seed generation—namely, the initial frame, bounding box trajectories, and even crash type category. In this setting, we sample 200 videos from the MM-AU validation set, and generate corresponding videos using their conditioning information. We do not directly compare generated videos to their exact ground-truth counterpart; instead, we treat these 200 ground-truth videos as a reference distribution that is semantically aligned with the generated one (e.g., similar vehicle types, scene layouts, and crash categories). FVD and JEDi are then computed between these two condition-aligned distributions. This setup ensures that the real and generated samples share similar structural priors, making the distributional comparison more meaningful. This method is used for FVD and JEDi results reported in Table 2 and Table 3 of the main paper.
@@
    Random GT Evaluation: Some baselines, such as AVD2, only provide generated videos without access to the original conditioning inputs, preventing us from constructing a semantically aligned ground-truth distribution. To ensure fair comparison in such cases, we instead sample 500 videos randomly from the MM-AU validation set to represent the ground-truth distribution. We then compare this to 200 generated samples from each model (including AVD2, Ctrl-Crash, Ctrl-V, and SVD Base). While this evaluation provides less precise alignment between real and generated content, it allows for inclusion of baselines where seed information is unavailable. This method is used for FVD and JEDi results reported in Table 1 of the main paper.
By using both evaluation protocols, we balance semantic alignment (when possible) with broad comparability, enabling fair and informative assessments across different classes of models.
In addition, we report frame-level metrics that directly assess perceptual quality relative to ground truth frames. These include LPIPS (Learned Perceptual Image Patch Similarity) [33], SSIM (Structural Similarity Index), and PSNR (Peak Signal-to-Noise Ratio). These metrics quantify visual similarity, structural consistency, and reconstruction fidelity, respectively, offering finer-grained insight into how well the generated frames preserve appearance and local coherence. To extend these frame-level metrics to video, we average the scores for all the frames of a given video.